OpenStudy (anonymous):

2√50 - 3√8 please simplify

OpenStudy (zzr0ck3r):

2sqrt50 = 5*2sqrt(2) 3sqrt(8) = 2*3sqrt(2) 10sqrt(2) - 6sqrt(2) = 4sqrt(2)

OpenStudy (zzr0ck3r):

2sqrt(5) = 2sqrt(25*2) = 2sqrt(2)sqrt(25) = 5*2*sqrt(2) the second line is done by the same reasoning

OpenStudy (anonymous):

okay, thanks!